OpenStudy (anonymous):

Carroll High School has a total of 94 boys and girls who play sports. If the number of boys is 9 more than four times the number of girls, how many boys play sports at this high school? 17 23 77 92 77 92

OpenStudy (anonymous):

say a1= number of boys a2= number of girls a1=\[4*a2+9\] a1+a2 =94 \[5*a2+9 = 94\] a2=\[85/5\] a2 =17 (number of girls) a1 = 77 (number of boys)